All disks started at zero GB.

hb: FGI-0004

ke: USN-0069

yg: USN-0023

Nothing to report except experiment started OK at 18:30.

Notes on tsys:

  • S-band reading high on ke all night.
  • X-band reading low on yg all night.
  • S-band reading high on hb all night. (e.g. ch09 – ch12 127,118,112,129 @ 2014.220.18:04:31.08)
  • Temperature values noted intermittently in additional notes of checklist.